OpenStudy (anonymous):

How many degrees are in 3pi/2 radians?

OpenStudy (zehanz):

A full circle is 360 degrees. If you take the Unit circle, the circumference is 2*pi*r=2*pi*1=2pi. Because the radian is the same unit as the length of the unit circle, there are also 2pi radians in a full circle. So remember: 360 degrees = 2pi radians. Now can you see the number of degrees in 3pi/2 (= 3/2*pi) radians?

OpenStudy (anonymous):

thx

OpenStudy (anonymous):

so whats the answer

hero (hero):

pi = 180 3pi/2 = 3(180)/2 = 3(90) = ?

OpenStudy (anonymous):

still confusion?

hero (hero):

You can't make it any more obvious than that.

OpenStudy (anonymous):

ya i get it thx

hero (hero):

So...what is it then?

OpenStudy (anonymous):

270

OpenStudy (anonymous):

What is the exact value of cos 45° as found on the unit circle?

hero (hero):

I'm only going to show you this once...I hope you understand it

OpenStudy (anonymous):

k

hero (hero):

\[\sin(0,30,45,60,90) = \frac{\sqrt{0,1,2,3,4}}{2}\] \[\cos(0,30,45,60,90) = \frac{\sqrt{4,3,2,1,0}}{2}\]